OpenStudy (anonymous):

find absolute extreme values of the function on the interval . f(x)=7(x)1/3-5x, [-3,4]. note: The first x is the power of 1/3

OpenStudy (anonymous):

are you allowed to use calculus?

OpenStudy (anonymous):

yes of course. you can use any level of calculus, the education system here in india is a bit ambiguous.

OpenStudy (anonymous):

well u only need to know derivatives.

OpenStudy (anonymous):

hmmm, i'm very confused about putting the interval values in the function.

OpenStudy (anonymous):

if i remember right the way to find the extremes is by f'(x)=0 On your interval the [ and ] mean that -3 and 4 may be included

OpenStudy (anonymous):

yeah, so after all that \[x ^{2/3}=7\div15\] \[x ^{2}=(7\div15)^3\] \[x = \pm7/15\sqrt{7/15}\] is it correct?